# Repeating partial disruptions and two-body relaxation

### v1

@@ Appendix A Expansion due to tidal oscillations
We report here a set of simulations that allow for the conversion of energy stored in oscillations into internal energy of the star. This is implemented through an expansion of the star, computed between step 3 and step 4 of the procedure outlined in 3.2. This is expected to happen only when the oscillations of the stellar structure are strongly non-linear. To estimate when this happens, we compare the total energy stored in oscillations with the internal energy of the star. For a polytrope model with index $n$ (Chandrasekhar(1957))
@@
In Fig. 7 and 8 we report the orbits at the threshold value of $r_{\mathrm{p}}$ and the overview of the outcomes at $q=0.1$, $q=0.01$ and $q=0.001$.

### v2

@@ Appendix A Expansion due to tidal oscillations
We report here a set of simulations that allow for the conversion of energy stored in oscillations into internal energy of the star. This is implemented through an expansion of the star, computed between step 3 and step 4 of the procedure outlined in 3.2. This is expected to happen only when the oscillations of the stellar structure are strongly non-linear. To estimate when this happens, we compare the total energy stored in oscillations with the internal energy of the star. For a polytrope model with index $n$ (Chandrasekhar, 1957)
@@
In Fig. 8 and 9 we report the orbits at the threshold value of $r_{\mathrm{p}}$ and the overview of the outcomes at $q=0.1$, $q=0.01$ and $q=0.001$.